# Non-Euclidean Erdős–Anning Theorems

### 12 Jan 2024 v1

@@ Abstract
The Erdős–Anning theorem states that every point set in the Euclidean plane with integer distances must be either collinear or finite. More strongly, for any (non-degenerate) triangle of diameter $\delta$, at most $O(\delta^{2})$ points can have integer distances from all three triangle vertices. We prove the same results for any strictly convex distance function on the plane, and analogous results for every two-dimensional complete Riemannian manifold of bounded genus and for geodesic distance on the boundary of every three-dimensional Euclidean convex set. Our proofs are based on the properties of additively weighted Voronoi diagrams of these distances.
@@ 1 Introduction
Our proofs replace the intersection properties of low-degree algebraic curves with the geometric and topological properties of additively weighted Voronoi diagrams.
@@ 3.1 Examples
For any $n$, one may form a complete Riemannian manifold from $n$ spheres, connected by cutting out disks from pairs of spheres, attaching a cylindrical tube between the boundaries of the two disks, and smoothing the parts of the surface where the disks meet the tubes. This construction may be performed abstractly as a topological space; it is not necessary to embed the spheres and tubes into three-dimensional space, allowing the lengths of the tubes to be adjusted. Let $S$ consist of one point per sphere, for spheres small enough that all distances within one sphere are less than $1/3$; then by adjusting the lengths of all the tubes, all distances between points of $S$ can be made equal to one. This construction produces arbitrarily large finite sets of points at unit distance from each other, with genus $O(n^{2})$. Because they all have the same distance, no three of these points belong to a shortest curve. By choosing the widths of the tubes from a quickly-decreasing sequence, it is possible to attach a countably-infinite number of tubes to each sphere, using the same construction to produce an infinite set of points at unit distance from each other, on a surface of infinite genus.
@@ 5 Discussion
We have shown that integer-distance point sets are finite or collinear in three broad classes of metric spaces each generalizing Euclidean distance in the plane. The next natural direction to look for a further extension of these results is in higher dimensions. Erdős and Anning state without proof that their results extend to $d$-dimensional Euclidean space [1]. However, the obvious attempt to generalize our proof to higher dimensions, by using intersection patterns of $(d+1)$-tuples of Voronoi cells, does not work for convex distance functions: there exist three-dimensional convex distance functions for which some 4-tuple of points is equidistant from arbitrarily many other points [15]. For Riemannian manifolds, the theorem does not generalize: one can smoothly distort the distances of Euclidean 3-space to allow countable sets of points to have integer distances without lining up along geodesics. Nevertheless, it would be interesting to study similar questions for higher-dimensional uniform metrics, such as hyperbolic space. It would also be of interest to strengthen the bound on diameter in Theorem 6 and Theorem 9, if possible.

### 22 Jan 2024 v2

@@ Abstract
The Erdős–Anning theorem states that every point set in the Euclidean plane with integer distances must be either collinear or finite. More strongly, for any (non-degenerate) triangle of diameter $\delta$, at most $O(\delta^{2})$ points can have integer distances from all three triangle vertices. We prove the same results for any strictly convex distance function on the plane, and analogous results for every two-dimensional complete Riemannian manifold of bounded genus and for geodesic distance on the boundary of every three-dimensional Euclidean convex set. As a consequence, we resolve a 1983 question of Richard Guy on the equilateral dimension of Riemannian manifolds. Our proofs are based on the properties of additively weighted Voronoi diagrams of these distances.
@@ 1 Introduction
Our proofs replace the intersection properties of low-degree algebraic curves with the geometric and topological properties of additively weighted Voronoi diagrams. Our results for Riemannian manifolds resolve a question posed in 1983 by Richard Guy [13], on the equilateral dimension of Riemannian manifolds, the maximum number of points at unit distance from each other in these manifolds. For complete Riemannian 2-manifolds of bounded genus, we bound the equilateral dimension by a function of the genus (Corollary 20) but for locally Euclidean incomplete Riemannian 2-manifolds, for complete Riemannian 2-manifolds of unbounded genus, and for complete Riemannian metrics on $\mathbb{R}^{3}$, the equilateral dimension is unbounded (Example 6 and Example 7).
@@ 3.1 Examples
For any $n$, one may form a complete Riemannian manifold from $n$ spheres, connected by cutting out disks from pairs of spheres, attaching a cylindrical tube between the boundaries of the two disks, and smoothing the parts of the surface where the disks meet the tubes. This construction may be performed abstractly as a topological space; it is not necessary to embed the spheres and tubes into three-dimensional space, allowing the lengths of the tubes to be adjusted. Let $S$ consist of one point per sphere, for spheres small enough that all distances within one sphere are less than $1/3$; then by adjusting the lengths of all the tubes, all distances between points of $S$ can be made equal to one. This construction produces arbitrarily large finite sets of points at unit distance from each other, with genus $O(n^{2})$. Because they all have the same distance, no three of these points belong to a shortest curve.
The connected sum of an infinite sequence of examples of this type for $n=4,5,6,\dots$, obtained by using tubes to connect the example for each $n$ to the examples for $n-1$ and $n+1$, forms a single complete Riemannian manifold of infinite genus in which one can find arbitrarily large (but nevertheless finite) sets of points that are all at unit distance from each other.
Although the topological structure of this example is complicated, it can be embedded in a distance-preserving way into a Riemannian metric on $\mathbb{R}^{3}$. For each $n$, choose the spheres and tubes of the $n$-point example to have a small enough radius that they can all be smoothly embedded within a unit ball of $\mathbb{R}^{3}$, coiling the tubes to cause them to have the desired lengths. Form the connected sum of these examples as before and embed it smoothly in $\mathbb{R}^{3}$ using disjoint unit balls for each term in the connected sum. Thicken the resulting smoothly embedded surface $S$ by a small amount (varying smoothly over the surface) chosen to be small enough that the thickened set remains smoothly embedded in $\mathbb{R}^{3}$, and give the resulting thickened set a Riemannian metric that agrees with the metric on $S$ (and the Euclidean distance of $\mathbb{R}^{3}$) in directions parallel to $S$, but that is longer in the direction perpendicular to $S$ by a factor that varies smoothly from very large for points on $S$ to one at the boundary of this thickened set. Then, for the remaining points of $\mathbb{R}^{3}$, outside the thickened set, use the usual Euclidean distance. By choosing this perpendicular lengthening factor large enough, it is possible to ensure that escaping from $S$ to the boundary of the thickened set takes more than unit distance, and that distances within $S$ are preserved.
Richard Guy asked whether the equilateral dimension of Riemannian manifolds, the maximum cardinality of a set of points in the manifold for which all distances are one, can be bounded by a function of the dimension of the manifold [13]. A positive answer was known only for the case of complete Riemannian manifolds of bounded curvature [22]. Example 6 provides a counterexample for locally Euclidean and topologically simple but incomplete Riemannian 2-manifolds. Example 7 provides a counterexample for complete Riemannian 2-manifolds of unbounded genus and for topologically simple complete Riemannian 3-manifolds. On the other hand, our generalization of the Erdős–Anning theorem to complete Riemannian 2-manifolds of bounded genus also provides a bound on the equilateral dimension of these manifolds, as a function of their genus (Corollary 20).
@@ 3.4 Erdős–Anning theorems for Riemannian manifolds
Corollary 20.
The equilateral dimension of a 2-dimensional complete Riemannian manifold of bounded genus $g$ is $O(g)$.
@@ 5 Discussion
We have shown that integer-distance point sets are finite or collinear in three broad classes of metric spaces each generalizing Euclidean distance in the plane. The next natural direction to look for a further extension of these results is in higher dimensions. Erdős and Anning state without proof that their results extend to $d$-dimensional Euclidean space [1]. However, the obvious attempt to generalize our proof to higher dimensions, by using intersection patterns of $(d+1)$-tuples of Voronoi cells, does not work for convex distance functions: there exist three-dimensional convex distance functions for which some 4-tuple of points is equidistant from arbitrarily many other points [16]. For Riemannian manifolds, the theorem does not generalize, as Example 7 shows. Nevertheless, it would be interesting to study similar questions for higher-dimensional uniform metrics, such as hyperbolic space. It would also be of interest to strengthen the bound on diameter in Theorem 6 and Theorem 9, if possible.